Find the value of
93
Find the value of
93
Find the value of
93

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

The above value can be written as,
= 9 × 9 x 9
​= 729

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
The value of the following number can be calculated as
9³ = 9×9×9 = 729
 
Hope it helps!
Thank you ✌️